OpenStudy (anonymous):

If 1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = y, then x = ?

OpenStudy (phi):

you can substitute y into the Left-hand side to get (1/2)x + (1/2)y= y which simplifies to x=y so if y=0, x=0. if y=1, x=1 and so on...

OpenStudy (anonymous):

\[1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = x*\sum_{k=1}^{+\infty}1/2^k\]

OpenStudy (anonymous):

1/2^k est une série de Rieman qui converge vers 1

OpenStudy (anonymous):

donc y=x*1 d'où x=y

OpenStudy (anonymous):

\[1/2+1/4+1/8+...+1/2^n=1-1/2^n\]qui tend vers 1 lorsque n tend vers +infinity

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome sorry i wrote it in french, i just realized you can use google translate, sorry again

OpenStudy (anonymous):

I know, my nick name makes you think I am french, right?

OpenStudy (anonymous):

hhh kind of yes, i originally speak arabic since i am from morocco

OpenStudy (anonymous):

haha :].